Feed aggregator
COMMENT ON New Bunch-Length System Installation
COMMENT ON Install/Connect Injector Girder 0L00 and Girder 0L01 Beamline to NEW Injector 1/4/Booster Cryomodule
Magnet Measurement:
We completed torque checks on QJ quad on girder just down beam of 1/4 Cryomodule. We also wired the QJ-6, hooked up LCW and turned LCW ON. Leak check is completed and in good order.
Matt-the Haimson correctors are ready to mount, if you need assistance just shout.
iocin1 restored
COMMENT ON ZA dipoles return LCW water line replacement (7 magnets) - SAD 2023
COMMENT ON Magnet warm region girder work - SAD 2023
Magnet Measurement:
We completed torque checks on QJ quad on girder just down beam of 1/4/ Cryomodule. We also wired the QJ-6, hooked up LCW and turned LCW ON. Leak check is completed and in good order.
Matt-the Hamison correctors are ready to mount, if you need assistance just shout.
COMMENT ON Spring SAD 2023 RF Work Tracking
COMMENT ON PSS system upgrade SAD 2023
Corrected field wiring of MBO dipole monitoring system (analog signals).
Corrected wiring of the East Arc Beam Stopper (disconnection of old, unused pressure switch cables).
Corrected safety relay feedback signal wiring of SL RF14 interface chassis.
Completed assembly of position switches on Aperture A4.
Laser stage measurement
Looked at beam spot on "virtual cathode" camera as a function of laser-stage position. 5 points per axis, not great but a useful first data set. Tad bit tedious because the camera software is about as crappy as that for the UITF camera, if not worse. Per usual, the only function the user needs is not there.
Figure 1: Center and width of Gaussian fits in x and y applied to beam profiles integrated over the other axis, respectively.
Figure 2: Resulting stage calibration.
Disregarding the sign inversion of y (sloppy analysis), we get about 0.97 mm per 1000 stage units in x and 0.91 in y.
iocsoftin00srfvac restored
COMMENT ON Spring SAD 2023 RF Work Tracking
Tried re-installing waveguide pieces in 1L26 but only two cavities, 4 and 8, were aligned close enough to complete installation. Other six cavities will require new pieces measured, cut and fitted. These cavities might be fitted with short waveguide adapter plates we have on hand but will most likely need to be broken apart and fitted depending on the offset we need to make it a closer fit. This will also require more of our ladder platforms, ratchets and multiple techs to hold up pieces.
injector work summary 6/1/23
Greg activated the NEGs inside the dif pump station downstream of the booster. Combined ion pump current 15uA and falling.
Marcy 'conditioned' the three NEG pumps in the chopper region. She will perform a full activation tomorrow. Then we can remove the pump cart to provide better access to chopper for HCO
Larry and Sam made good progress with waveguides to booster
S&A completed 2B alignment of the phase2 girders, nice.
Phil provided assistance to SSG, the SSG boxes are now bolted to floor near beamline/A3/A4 and MBO dipole
COMMENT ON Modify 0L02 RF drive system for booster upgrade
Waveguide installation for the Booster has started with the 400W waveguide (cavity 7) being test fitted. All gaskets will be installed and flanges properly torqued tomorrow. Klystron waveguide (cavity 8) will be also worked on tomorrow. Techs: L. Farrish and S. Holben
COMMENT ON Box Power Supply Preventative Maintenance and Interlock Checks Spring 2023
NEG activations: 0I02, 0I04 and 0I05
COMPLETED VDP0L01 Cleaning/Vacuuming Interior Base Plate and VIP Change-out
COMMENT ON Dogleg Power Supply PM's
COMMENT ON Phase 2 Injector Magnets Cables Pull and Termination for New Magnets to be Installed
iocsoftinj1 restored
COMMENT ON Re-connect/Install West ARC Beamline
Pages
